# Holographic Tensor Networks as Tessellations of Geometry

### v2

@@ Abstract
Holographic tensor networks serve as toy models for the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence, capturing many of its essential features in a concrete manner. However, existing holographic tensor network models remain far from a complete theory of quantum gravity. A key obstacle is their discrete structure, which only approximates the semi-classical geometry of gravity in a qualitative sense. In Lin:2024dho , it was shown that a network of partial-entanglement-entropy (PEE) threads, which are bulk geodesics with a specific density distribution, generates a perfect tessellation of AdS space. Moreover, such PEE-network tessellations can be constructed for more general geometries using the Crofton formula. In this paper, we assign a quantum state to each vertex in the PEE network and develop two holographic tensor network models: the factorized PEE tensor network, which takes the form of a tensor product of EPR pairs, and the random PEE tensor network. In both models we reproduce the exact Ryu–Takayanagi formula by showing that the minimal number of cuts along a homologous surface in the network exactly computes the area of this surface.
@@ 1 Introduction
The tensor network approach Swingle:2009bg ; Swingle:2012wq ; Qi:2013caa ; Miyaji:2015fia ; Czech:2015kbp ; Harlow:2016vwg ; Vasseur:2018gfy ; Dong:2018seb ; Hung:2019zsk ; Bao:2019fpq provides a concrete framework to explore the structure of the AdS/CFT correspondence 111see Pastawski:2015qua ; Hayden:2016cfa ; Yang:2015uoa ; Qi:2013caa ; Bao:2019fpq ; Cheng:2022ori ; Dong:2023kyr for an incomplete list of recent progress.. More explicitly, we start with a space filled with tensors representing quantum states on each site in the space, then we contract the “legs” (indices) to connect adjacent tensors thus get a tensor network, which plays the role of the background geometry. The uncontracted legs of tensors in the bulk and the boundary are regarded as the bulk and the boundary degrees of freedom respectively, and the tensor network provides a quantum-error-correcting code that encodes the bulk states into the boundary Hilbert space Dong:2016eik ; Hayden:2016cfa . Many toy models of tensor network to simulate the AdS/CFT correspondence have been proposed (see also Chandra:2023dgq ; Chen:2024unp ; Bao:2024ixc ; Kaya:2025vof ; Akella:2025owv for recent interesting developments), based on different structures of the networks and different quantum states associated to the tensors. One important holographic feature for these toy models is that, the entanglement entropy is given by the minimal number of cuts along a homologous surface in the network, which matches the area of the minimal surface of the celebrated Ryu-Takayanagi (RT) formula Ryu:2006bv ; Hubeny:2007xt ; Casini_2011 ; Lewkowycz:2013nqa ; Faulkner:2013ana ; Engelhardt:2014gca .
@@
Nevertheless, in most of these toy models their structures of the tensor network are usually discrete and only capture the features of the background geometry in a qualitative way. There is a big gap between number of cuts in a discrete network and the area of a surface in a smooth Riemannian manifold. Recently, a continuous network as a perfect tessellation of the background AdS geometry was constructed using the so-called partial-entanglement-entropy (PEE) threads Lin:2023rxc ; Lin:2024dho , which are bulk geodesics with a particular density distribution determined by the PEE structure of the boundary CFT. In such a network, the number of cuts along a surface exactly reproduces the area of this surface. In this work, we will develop two toy models of tensor network based on the PEE threads (or the PEE tensor network for short) in $d$-dimensional Poincaré AdS space. The first one is the factorized PEE tensor network constructed via tensor product of EPR pairs. It enjoys similar properties to the Harlow-Pastawski-Preskill-Yoshida (HaPPY) code Pastawski:2015qua which is built from perfect tensors, and perfectly reproduces the RT formula for spherical regions in the vacuum CFT state. In the second model, following Hayden:2016cfa ; Vasseur:2018gfy ; Jia:2020etj ; Cheng:2022ori ; Dong:2023kyr ; Kaya:2025vof we associate a random state to each bulk site to build the random PEE tensor network, where we can exactly reproduce the RT formula for generic boundary regions.
@@ 2 The network of PEE threads
Here the regions are all non-overlapping, $A\cup\bar{A}$ makes a pure state and $S_{A}$ denotes the entanglement entropy of $A$. The solutions for these requirements were first studied in Casini_2011 as the mutual information that is additive in some special theories, and then the solutions are understood as the PEE in Wen:2019iyq , which is in general not a mutual information. Later it was gradually realized that, the requirement of normalization should not be imposed for an arbitrary region. In the case of the vacuum CFT state on a plane, the solution only exists if we impose the normalization requirement to spherical regions Wen:2019iyq . See also Vidal:2014aal ; Wen:2018whg ; Wen:2018mev ; Wen:2019iyq ; Kudler-Flam:2019oru for other prescriptions to construct the PEE which satisfy the above requirements, and see Basu:2023wmv ; Rolph:2021nan ; Lin:2022aqf ; Lin:2023orb ; Wen:2021qgx ; Wen:2022jxr ; Wen:2024uwr ; Camargo:2022mme for related discussions on PEE.
@@
In the context of AdS/CFT, the vacuum CFTd state on a plane is dual to the Poincaré AdSd+1 space. The two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ was also proposed to have a natural holographic dual Lin:2023rxc ; Lin:2024dho , which is the density of the bulk geodesics connecting the boundary sites $\mathbf{x}$ and $\mathbf{y}$. The bulk geodesics whose density are determined in such a way are denoted as the PEE threads. Let us set that each PEE thread gives unit contribution to the PEE, then (1) captures the number of PEE threads connecting the two boundary regions $A$ and $B$. The collection of all the PEE threads forms a continuous, dense network permeating the entire time slice of the Poincaré AdSd+1, which we call the PEE network (see an example of $d=2$ in Fig.1). Remarkably, it was found that Lin:2023rxc ; Lin:2024dho the PEE network is a tessellation of the static AdS space in the following sense:
@@
Then for any bulk surface $\Sigma$, the area can be reproduced by counting the number $\mathcal{N}\left(\Sigma\right)$ of intersections between $\Sigma$ and the PEE network,
@@
Here $\mathcal{M}$ is a constant time slice in the Poincaré AdSd+1, while $\partial\mathcal{M}$ is the boundary where the vacuum CFT state lives, and $\omega_{\Sigma}\left(\mathbf{x},\mathbf{y}\right)$ represents of the number of intersections between $\Sigma$ and the particular PEE thread connecting boundary sites $\mathbf{x}$ and $\mathbf{y}$.
@@
Here, the integrand $\left(\Gamma\cap\Sigma\right)$ represents the number of intersections between any geodesic $\Gamma$ in $\text{Geod}\left(M\right)$ and $\Sigma$, and the Kinematic measure $d\Gamma$ specifies the density of geodesics in some parameter space. For the case of a time slice $\mathcal{M}$ in the AdSd+1, all the geodesics can be parameterized by its boundary endpoints $\left(\mathbf{x},\mathbf{y}\right)$, and the Kinematic measure $d\Gamma$ is given by,
@@
where $\ell\left(\mathbf{x},\mathbf{y}\right)$ is the geodesic length of the geodesic connecting $\mathbf{x}$ and $\mathbf{y}$. Notably, the above Kinematic measure for AdS space is proportional to $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ given in (2), hence we can check that (4) exactly aligns with (3).
One profound implication of the above prescription is that, minimizing the area of a surface $\Sigma_{A}$ homologous to any boundary region $A$ is equivalent to minimizing the number of intersections between $\Sigma_{A}$ (see the red line in Fig.1) and the PEE network. This further implies that we can exactly reproduce the Ryu-Takayanagi (RT) formula for holographic entanglement entropy by developing proper tensor network models where the entanglement entropy of a boundary region is captured by the minimal number of intersections between a homologous surface and the PEE network Lin:2023rxc ; Lin:2024dho ; Wen:2024uwr . Briefly speaking, the RT formula for the boundary region $A$ can be reformulated as follows,
@@
Figure 1: Extracted from Wen:2024uwr . The PEE network as a tessellation of a time slice of AdS3. Here the blue lines represent the PEE threads, the red line is a surface $\Sigma_{A}$ homologous to the boundary region AA.
@@ 3 Geometric setup for bulk tensors
In a model of tensor network, a vertex at site $\mathbf{z}$ with a set of legs corresponds to a quantum state represented by a local tensor $\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})$, where the index $u_{i}$ characterizes the leg that is associated to a Hilbert space $\mathcal{H}_{i}$ spanned by the basis vectors $\{|u_{i}\rangle\}$. Then the quantum state in the Hilbert space $\bigotimes_{i}\mathcal{H}_{i}$ settled at $\mathbf{z}$ is defined by (assuming Einstein summation hereafter),
$|\mathcal{T}(\mathbf{z})\rangle\equiv\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})|u_{1}\rangle|u_{2}\rangle\cdots.$ (7)
@@
Now we develop the PEE tensor networks. Consider any site $\mathbf{z}$ in a time slice $\mathcal{M}$ of the bulk space, for any PEE thread emanating from it to the boundary, we assign one leg that lies along this PEE thread, hence build a one-to-one correspondence between the PEE threads and the legs connecting to $\mathbf{z}$ (see Fig.2). In the case of Poincaré AdS, any PEE thread emanating from $\mathbf{z}$ intersects the boundary with a unique boundary site $\mathbf{x}$. We denote this thread as $C_{\mathbf{z}}(\mathbf{x})$, and $\mathbf{x}$ can be used to parameterize all the threads or legs associated to $\mathbf{z}$. Let us denote the Hilbert space and basis vectors on each leg as $\mathcal{H}_{\mathbf{x}}$ and $|k_{\mathbf{x}}\rangle$, such that the state (7) is rewritten as
$|\mathcal{T}(\mathbf{z})\rangle\equiv\mathcal{T}_{\mathbf{k}}(\mathbf{z})\left(\bigotimes_{\mathbf{x}\in bdy}|k_{\mathbf{x}}\rangle\right),$ (8)
@@
where $|\mathbf{z}\mathbf{z^{\prime}}\rangle=\frac{1}{\sqrt{v}}\sum_{k=1}^{k=v}|k\rangle_{\mathbf{z}}|k\rangle_{\mathbf{z^{\prime}}}$ is a maximally entangled state defined on any pair of legs connecting the nearby sites $\mathbf{z}$ and $\mathbf{z^{\prime}}$. These contractions connect all the legs into complete PEE threads (see Fig.3) and eventually we get the PEE tensor networks.

### v3

@@ Abstract
Holographic tensor networks serve as toy models for the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence, capturing many of its essential features in a concrete manner. However, existing holographic tensor network models remain far from a complete theory of quantum gravity. A key obstacle is their discrete structure, which only approximates the semi-classical geometry of gravity in a qualitative sense. In Lin:2024dho , it was shown that a network of partial-entanglement-entropy (PEE) threads, which are bulk geodesics with a specific density distribution, generates a perfect tessellation of AdS space. Moreover, such PEE-network tessellations can be constructed for more highly symmetric geometries using the Crofton formula. In this paper, we assign a quantum state to each vertex in the PEE network and develop several holographic tensor network models: (1) the factorized PEE tensor network, which takes the form of a tensor product of EPR pairs; (2) the HaPPY-like PEE tensor network constructed from perfect tensors; and (3) the random PEE tensor network. In all these models, we reproduce the exact Ryu-Takayanagi formula by showing that the minimal number of cuts along a homologous surface in the network exactly equals the area of that surface.
@@ 1 Introduction
The tensor network approach Swingle:2009bg ; Swingle:2012wq ; Qi:2013caa ; Miyaji:2015fia ; Czech:2015kbp ; Harlow:2016vwg ; Vasseur:2018gfy ; Dong:2018seb ; Hung:2019zsk ; Bao:2019fpq provides a concrete framework to explore the structure of the AdS/CFT correspondence 111see Pastawski:2015qua ; Hayden:2016cfa ; Yang:2015uoa ; Qi:2013caa ; Bao:2019fpq ; Cheng:2022ori ; Dong:2023kyr for an incomplete list of recent progress.. More explicitly, we start with a space filled with tensors representing quantum states on each site in the space, then we contract the “legs” (indices) to connect adjacent tensors thus get a tensor network, which plays the role of the background geometry. The uncontracted legs of tensors in the bulk and the boundary are regarded as the bulk and the boundary degrees of freedom respectively, and the tensor network provides a quantum-error-correcting code that encodes the bulk states into the boundary Hilbert space Dong:2016eik ; Hayden:2016cfa . Many toy models of tensor network to simulate the AdS/CFT correspondence have been proposed (see also Chandra:2023dgq ; Chen:2024unp ; Geng:2025efs ; Bao:2024ixc ; Kaya:2025vof ; Akella:2025owv ; Akers:2021pvd ; Akers:2022zxr ; Akers:2024ixq ; Akers:2023obn ; Akers:2024wab ; Akers:2024pgq for recent interesting developments), based on different structures of the networks and different quantum states associated to the tensors. One important holographic feature for these toy models is that, the entanglement entropy is given by the minimal number of cuts along a homologous surface in the network, which matches the area of the minimal surface of the celebrated Ryu-Takayanagi (RT) formula Ryu:2006bv ; Hubeny:2007xt ; Casini_2011 ; Lewkowycz:2013nqa ; Faulkner:2013ana ; Engelhardt:2014gca .
Nevertheless, in most of these toy models their structures of the tensor network are usually discrete and only capture the features of the background geometry in a qualitative way. There is a big gap between number of cuts in a discrete network and the area of a surface in a smooth Riemannian manifold. Note that the study of continuous tensor network (cTN) models is not new in the literature; see Verstraete:2010ft for the initial proposal and Haegeman:2012uk ; Jennings:2012bsr ; Tilloy:2018gvo ; Tilloy:2021yre for subsequent developments. Relevant applications in high energy physics can be found in Shachar:2021vbu and Vidal:2008zz ; Haegeman:2011uy ; Nozaki:2012zj . In these models, the continuum nature is achieved by taking the infinitesimal limit of the lattice spacing in the underlying lattice configurations. Despite significant progress in approximating genuine continuous quantum field theories, the information about the background geometry is not fully encoded in the tensor network states constructed in this way, leaving the continuum simulation of such theories far from complete.
@@
Recently, a continuous network as a perfect tessellation of the background AdS geometry was constructed using the so-called partial-entanglement-entropy (PEE) threads Lin:2023rxc ; Lin:2024dho , which are bulk geodesics with a particular density distribution determined by the PEE structure of the boundary CFT. In such a network, the number of cuts along a surface exactly reproduces the area of this surface. In this work, we will develop three toy models of tensor network based on the PEE threads (or the PEE tensor network for short) in $d$-dimensional Poincaré AdS space. The first one is the factorized PEE tensor network constructed via tensor product of EPR pairs. The second one is a HaPPY-like (Harlow-Pastawski-Preskill-Yoshida, Pastawski:2015qua ) tensor network constructed from perfect tensors. It functions as a quantum error-correcting code and exactly reproduces the RT formula for connected regions in the vacuum CFT state. In the third model, following Hayden:2016cfa ; Vasseur:2018gfy ; Jia:2020etj ; Cheng:2022ori ; Dong:2023kyr ; Kaya:2025vof we associate a random state to each bulk site to build the random PEE tensor network, where we can exactly reproduce the RT formula for generic boundary regions.
The paper is organized as follows. In section 2, we give an introduction to the network of PEE threads. In section 3, we establish the geometric setup for our tensor network models. We introduce the factorized PEE tensor network in section 4 and give an example for the HaPPY-like PEE tensor network in section 5, and then introduce the random PEE tensor network in section 6. We end the paper with discussions in section 7 and give a geometric description for the factorized PEE tensor network in appendix A.
@@ 2 The network of PEE threads
Here the regions are all non-overlapping, $A\cup\bar{A}$ makes a pure state and $S_{A}$ denotes the entanglement entropy of $A$. The solutions for these requirements were first studied in Casini:2008wt as the mutual information that is additive in some special theories, and then the solutions are understood as the PEE in Wen:2019iyq , which is in general not a mutual information  [^2] Here, let us clarify the distinction between the PEE and EMI Casini:2008wt . In fact, the additivity together with the requirements 3-7 is already enough to determine the bilinear form (2) of both the EMI and the PEE, up to a constant coefficient. This bilinear structure exists in the vacuum state of a generic CFT. The key difference between the PEE and the EMI is that the EMI is a mutual information, and thus satisfies the normalization property as a consequence. The PEE, in contrast, is not a mutual information; the normalization property must be imposed as a requirement in order to define it. The study of EMI originated in the context of two-dimensional free massless fermions, where the mutual information was found to be additive. It was natural to assume that similar theories with additive mutual information might exist in higher dimensions, and these were termed EMI models. However, the existence of EMI models has been justified only for two-dimensional free massless fermions. Moreover, it was later shown that such models do not exist in higher-dimensional actual CFTs or in any of their limits Agon:2021zvp . See Bueno:2015rda ; Casini:2015woa ; Bueno:2019mex for the evaluation of entanglement entropies in the EMI models, obtained by taking limits in accordance with the normalization property. The concept of PEE, on the other hand, was motivated by a natural slicing of the entanglement wedge in holography Wen:2018whg , and by the additive entanglement structure conjectured to describe the spatial distribution of entanglement—namely, the entanglement contour function Vidal:2014aal . In one spatial dimension, a general proposal for the PEE Wen:2018whg ; Wen:2020ech exists, which defines the PEE as an additive linear combination of subsystem entanglement entropies; this was later referred to as the ALC proposal for PEE. In Wen:2019iyq , it was shown that the ALC proposal satisfies all seven requirements (including normalization) for generic theories in one spatial dimension, and that it is not a mutual information. It coincides with the EMI only in the special case of two-dimensional massless free fermions. For higher-dimensional theories, the bilinear form of the EMI could serve as an ideal candidate for the PEE, provided one abandons the assumption that the quantity represents a mutual information Wen:2019iyq . Even so, the bilinear structure determined by additivity and requirements 3-7 is fixed up to a constant coefficient. If the normalization requirement is imposed for all regions, then no solution exists in general theories that satisfies all seven requirements. In our work, we are particularly interested in the PEE in holographic CFTs, and we have recently found that the coefficient can be consistently determined by imposing the normalization requirement only for spherical regions, thereby ensuring the existence of a solution Lin:2023rxc ; Lin:2024dho . This new interpretation is precisely what makes the bilinear structure applicable to the vacuum state of holographic CFTs. . Later it was gradually realized that, the requirement of normalization should not be imposed for an arbitrary region. In the case of the vacuum CFT state on a plane, the solution only exists if we impose the normalization requirement to spherical regions Wen:2019iyq . See also Vidal:2014aal ; Wen:2018whg ; Wen:2018mev ; Wen:2019iyq ; Kudler-Flam:2019oru ; Wen:2024yny for other prescriptions to construct the PEE which satisfy the above requirements, and see Basu:2023wmv ; Rolph:2021nan ; Lin:2022aqf ; Lin:2023orb ; Wen:2021qgx ; Wen:2022jxr ; Wen:2024uwr ; Camargo:2022mme for related discussions on PEE.
@@
In the context of AdS/CFT, the vacuum CFTd state on a plane is dual to the Poincaré AdSd+1. The two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ was also proposed to have a natural holographic dual Lin:2023rxc ; Lin:2024dho , which is the density of the bulk geodesics connecting the boundary sites $\mathbf{x}$ and $\mathbf{y}$. The bulk geodesics whose density are determined in such a way are denoted as the PEE threads. These PEE threads can be described by vector fields, with the norm of each vector field capturing the density of threads. For instance, for each point on the AdS boundary, there exists a vector field that describes all the threads emanating from that point (see Fig.1a for illustration and see Lin:2023rxc for the explicit formula of the vector fields). The full partial entanglement entropy (PEE) network is then formed by the superposition of all such vector fields (see an example of $d=2$ in Fig.1b). In this context, the number of geodesics are countless but one can compute the “flux” associated with these vector fields. For example, in (1) the right hand side captures the flux of threads connecting boundary regions $A$ and $B$. Additionally, one can compute the total number of threads crossing through a given bulk surface $\Sigma$. It is important to note that the PEE threads are unoriented curves; each time a thread intersects $\Sigma$, it contributes positively to the total crossing count. Therefore, when computing this quantity, which differs from the conventional net flux in that contributions from opposite directions add rather than cancel, the threads should be weighted by the number of times they intersect $\Sigma$.
$\mathbf{x}$
Figure 1: In the left figure, the black solid line is the AdS boundary, and the dashed lines are geodesics or PEE threads in AdS space. The distribution of the PEE threads emanating from a boundary site $\mathbf{x}$ can be described by a vector field $V_{\mathbf{x}}^{\mu}$, whose integral curves coincide with the PEE threads and whose norm represents the density of these threads. The right figure, taken from Wen:2024uwr , shows the PEE network as the superposition of all vector fields $V_{\mathbf{x}}^{\mu}$, resulting in a tessellation of a time slice of AdS3. In the figure, the blue lines represent the PEE threads, and the red line is a surface $\Sigma_{A}$ homologous to the boundary region $A$.
In discrete networks, concepts such as the number of legs or the number of intersections between a surface and the network are commonly used. We can adopt this language, hence techniques in the discrete models can be used, by defining how much each thread contributes to the flux, or the total crossing each time it passes through $\Sigma$. If we set the Hilbert space dimension on each leg to be $v$, then the contribution from each PEE thread to the flux is simply $\log v$. Here we use $\log v$ instead of $\log_{2}v$ to adhere to the convention in high energy physics. Note that the flux or the total crossing, is the essential physical quantity, rather than the number of threads or intersections which depends on our choice of $v$. We set $\log v=1$ for the later discussions for convenience, then the flux exactly captures the number of the PEE threads, while the total crossing captures the number of intersections.
Remarkably, it was found that Lin:2024dho the PEE network is a perfect tessellation of the static AdS space in the following sense:
@@
In the above statement, we used $\log v=1$ to replace the total crossing with the number of intersections. But this is not essential: any other choice of $\log v$ merely rescales the number or the density of intersections by an overall factor $1/\log v$. Then for any bulk surface $\Sigma$, the area can be reproduced by counting the number $\mathcal{N}\left(\Sigma\right)$ of intersections between $\Sigma$ and the PEE network,
@@
[…]
@@
Here, the integrand $\left(\Gamma\cap\Sigma\right)$ represents the number of intersections between any geodesic $\Gamma$ in $\text{Geod}\left(M\right)$ and $\Sigma$, and the Kinematic measure $d\Gamma$ specifies the density of geodesics in some parameter space. For the case of a time slice $\mathcal{M}$ in the AdSd+1, all the geodesics can be parameterized by its boundary endpoints $\left(\mathbf{x},\mathbf{y}\right)$, and the Kinematic measure $d\Gamma$ is given by Czech:2016xec ; Huang:2015xca ; Lin:2024dho ,
@@
where $\ell\left(\mathbf{x},\mathbf{y}\right)$ is the geodesic length of the geodesic connecting $\mathbf{x}$ and $\mathbf{y}$. Notably, the above Kinematic measure for AdS space is proportional to $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ given in (2), hence we can check that (4) exactly aligns with (3). Note that for a generic manifold, the isometry group may be small, so there is no canonical “invariant” measure on the space of geodesics that is independent of location. The elegant formula (4) applies only to highly symmetric manifolds such as AdS space, Euclidean space, or the sphere; its generalization to more generic manifolds is highly nontrivial.
One profound implication of the above prescription is that, minimizing the area of a surface $\Sigma_{A}$ homologous to any boundary region $A$ is equivalent to minimizing the number of intersections between $\Sigma_{A}$ (see the red line in Fig.1b) and the PEE network. This further implies that we can exactly reproduce the Ryu-Takayanagi (RT) formula for holographic entanglement entropy by developing proper tensor network models where the entanglement entropy of a boundary region is captured by the minimal number of intersections between a homologous surface and the PEE network Lin:2023rxc ; Lin:2024dho ; Wen:2024uwr . Briefly speaking, the RT formula for the boundary region $A$ can be reformulated as follows,
@@ 3 Geometric setup for bulk tensors
In a model of tensor network, a vertex at site $\mathbf{z}$ with a set of legs corresponds to a quantum state represented by a local tensor $\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})$, where the index $u_{i}$ characterizes the leg that is associated to a Hilbert space $\mathcal{H}_{i}$ spanned by the basis vectors $\{|u_{i}\rangle_{\mathbf{z}}\}$. Then the quantum state in the Hilbert space $\bigotimes_{i}\mathcal{H}_{i}$ settled at $\mathbf{z}$ is defined by (assuming Einstein summation hereafter),
$|\mathcal{T}(\mathbf{z})\rangle\equiv\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})|u_{1}\rangle_{\mathbf{z}}|u_{2}\rangle_{\mathbf{z}}\cdots.$ (7)
@@
Now we develop the PEE tensor networks. Consider any site $\mathbf{z}$ in a time slice $\mathcal{M}$ of the bulk space, for any PEE thread emanating from it to the boundary, we assign one leg that lies along this PEE thread, hence build a one-to-one correspondence between the PEE threads and the legs connecting to $\mathbf{z}$ (see Fig.2). In the case of Poincaré AdS, any PEE thread emanating from $\mathbf{z}$ intersects the boundary with a unique boundary site $\mathbf{x}$. We denote this thread as $C_{\mathbf{z}}(\mathbf{x})$, and $\mathbf{x}$ can be used to parameterize all the threads or legs associated to $\mathbf{z}$. Let us denote the Hilbert space and basis vectors on each leg as $\mathcal{H}_{\mathbf{x}}$ and $|k_{\mathbf{x}}\rangle_{\mathbf{z}}$, such that the state (7) is rewritten as
$|\mathcal{T}(\mathbf{z})\rangle\equiv\mathcal{T}_{\mathbf{k}}(\mathbf{z})\left(\bigotimes_{\mathbf{x}\in bdy}|k_{\mathbf{x}}\rangle_{\mathbf{z}}\right),$ (8)
@@
where $|\mathbf{z}\mathbf{z^{\prime}}\rangle=\frac{1}{\sqrt{v}}\sum_{k=1}^{k=v}|k\rangle_{\mathbf{z}}|k\rangle_{\mathbf{z^{\prime}}}$ is a maximally entangled state defined on any pair of legs connecting the nearby sites $\mathbf{z}$ and $\mathbf{z^{\prime}}$. Such construction is also referred to as projected entangled pair states (PEPS)Verstraete:2003wvw ; Verstraete:2004cf ; Verstraete:2008cex ; Cirac:2020obd in condensed matter physics. This projection can be understood as the gluing of nearby geometries through the imposition of identical boundary conditions when performing the path integral in gravitational theories. These contractions connect all the legs into complete PEE threads (see Fig.3) and eventually we get the PEE tensor networks.